# Affordances and Design Principles of The Political Left and Right

### v1

@@ Abstract.
Like any form of technology, social media services embed values. To examine how societal values may be present in these systems, we focus on exploring political ideology as a value system. We organised four co-design workshops with political representatives from five major parties in Finland to investigate what values they would incorporate into social media services. The participants were divided into one right-leaning group, two left-leaning groups, and one mixed group. This approach allows us to examine how their political ideologies, i.e., value systems, influenced the design of social media. We analysed produced artefacts (early-stage paper mockups) to identify different features and affordances for each group and then contrasted the ideological compositions. Our results revealed a clear distinction between groups: right-leaning groups favoured market-based visibility, while left-leaning groups rejected such design principles in favour of open profile work. Additionally, we found tentative differences in design outcomes along the liberal–conservative dimension. These findings underscore the importance of acknowledging existing political value systems in the design of social computing systems. They also highlight the need for further research to map out political ideologies in technology design.
@@
copyright: $c$
@@ 1. Introduction
Researchers have often focused on end-users and less on decision-makers, such as political elites, when considering different values and their influence on design. At the same time, the government is a major provider of services, facilitated increasingly through digital means. Even more importantly, in democratic societies, the elected representatives have the legislative power – they establish laws increasingly related to technology, as shown in the EU AI Act or the California Consumer Privacy Act. Both involve value-based decisions about the rights and responsibilities of citizens and their expected roles. In the political system, there are reasons to suggest that such decisions mirror a wider set of beliefs and assumptions (or political ideologies), as well as political realities based on different balances of power among representatives. Members of Congress, Parliament or city councils, and other representatives rarely have been studied as stakeholders in the design of digital systems.
@@
We examined the outcomes of the workshops as the participants’ suggestions and analysed those towards the specific features and implicit affordances. Based on the literature on social media platforms (Boyd, 2010; Bucher and Helmond, 2017; DeVito et al., 2017; Van Raemdonck and Pierson, 2021), we categorised affordances and connected them to underlying values to uncover ideological tensions between the different design solutions.
We next discuss relevant literature to present political ideologies as values and discuss social media design further. Following this, we describe the organisation of the co-design workshops. In this paper, we focus only on the final artefacts of the co-design activity, which exemplify the value-based decisions the groups made during the design workshop. We present a method to analyse these artefacts and summarise observations from the artefacts. We observe that right-leaning groups supported market-based dynamics while left-leaning groups supported diverse self-presentation. Mixed groups seemed to have both kinds of features. Before concluding our work, we discuss the importance of ideological diversity in design activities. Our work makes the following contributions focused on the design of information technology:
@@
  • We show how political ideologies (right-left) shape social media design and identify design principles of flexible identity and competitive logic to be specific to the left- and right-leaning parties.
@@ 2. Background
Our work is based on the well-established observation that technology design and development are not value-neutral activities (among many others Nissenbaum, 2005; Winner, 1985). Winner (1985) infamously asked if artefacts have politics following a story about the bridges on Long Island, New York, USA. The bridges were constructed too low for public transportation buses to pass under them, thus removing access to the area for people using public transportation, primarily black people at that time. He suggests that this is related to the designers’ social class bias and racial prejudice. Therefore, the bridges became manifestations of their designer’s political values. In the context of human–computer interaction, Keyes et al. (2019) has similarly called out the often apolitical nature of work, which they say to be implicitly neo-liberal in nature, and thus calling out for more diverse ideological work.
@@ 2.1. Political Ideologies as a Set of Values
Across various disciplines, there are several ways to conceptualise values. Political scientists often focus on political ideologies instead of values. Political ideologies are a set of values, beliefs, and attitudes people share that impact how they perceive and act in the world (Denzau and North, 1994). Instead of single values, they formulate a value system. Within representative bodies, people with similar ideologies form a political party (or a social movement in their early stages) to collaborate in advocating their values. The existence of cohesive parties allows one to map out ideologies’ dimensions; in the United States of America, this is often represented as a single dimension (republican–democrat) where various societal conflicts, such as economic stances, the role of the government, or racial assumptions, are positioned.
@@
In Europe (where our study is situated), the multiparty system makes more differentiated ideologies apparent. Instead of a single axis, a two-dimensional model is used across these countries (see Figure 1). The first axis corresponds to different values in terms of the degree of economic freedom (left–right), and the second axis corresponds to the level of desirable societal diversity (conservative–liberal or green-alternative-liberal–traditional-authoritarian-nationalist) (CAUGHEY et al., 2019; Kriesi et al., 2006). For example, a supporter or member of a left-liberal party, such as social democrats or the greens, might push for the right to self-define gender identity and extensive government oversight in the economic operation. In contrast, a supporter or member of a right-conservative party, such as Christian democrats or conservative parties, might push for traditional gender roles and limit the government’s role in the economy. In general, while the model is simplistic, it has provided extensive value for empirical research on political decision-making, political opinions, and voting (e.g., CAUGHEY et al., 2019; Huber and Inglehart, 1995; Hooghe et al., 2002; Knight, 2006). There is mixed evidence that left-leaning political parties are adapting information and communication technology more extensively to allow contact and participation in party functions (González-Cacheda and Cancela Outeda, 2024).
@@
As ideologies are systems based on a set of values, these values can manifest themselves in the design and use of digital services. Examining the design of social media services, Grön and Nelimarkka (2020) illustrate that the left–right axis relates to different degrees of paternalistic values: left-leaning participants supported a higher degree of paternalism, while right-leaning participants highlighted the importance of individual responsibility. Their work highlights ideologies as an area of political competition, where different ideologies prefer design approaches aligned with their worldview. Similarly, Détienne et al. (2019) built on the concept of ideologies to understand community memberships in conflict. For example, they identified an ideological conflict in the editing of Wikipedia articles. They observed strong disagreement between authors with a religious point of view and those with a scientific point of view, leading to an edit war and personal attacks on the platform. This instance underscores the potential for conflict and disagreement between political ideologies, making it clear that finding a compromise on an agreed solution can be difficult and sometimes impossible. This work highlights the importance of reflecting more expansive sets of values (or ideologies) and understanding them through wider groups, even if their ideologies do not fit the two-dimensional political ideology model presented above.
@@
Although ideologies and their impact on information systems design have received less attention in human–computer interaction, the above examples illustrate the value of approaching not only individual values but sets of values. The focus on political ideologies – pushed forward by political parties – provides the additional benefit of reflecting the legislative and executive processes at the societal level.
@@ 2.2. Values in the Context of Social Media Services
Like any socio-technical arrangement, social media services incorporate and display values. The most prominent value tension in contemporary public discourse concerns social media companies’ for-profit operations and their contrast with other values, such as human rights. Thus, the discourse has a focus on capitalism. For example, Zuboff (2018) has argued that we are now in the era of surveillance capitalism, in which the collection and use of personal data has become a widespread business model. Social media companies use such personal data for targeted advertising, but while doing so, limit our right to privacy. An alternative line of criticism is the focus on user engagement and attention as a (capitalist) metric, leading to the promotion of affective content and thus having negative implications for political discourse on these platforms.
Given these concerns, there are proposals on alternative value bases for social media services. Jia et al. (2024) discuss how the news-feed algorithm could promote content more aligned with democratic ideals by sorting content so that it does not show support for undemocratic practices, partisan animosity or violence, or social distrust. Others have proposed different normative bases for rethinking social media services (for summary, see Goldberg, 2016). For example, the ideas that social media services could serve as a place for political deliberation have been repeated in the literature (Semaan et al., 2014, 2015, among others). These examples again illustrate the critical role of values as a standing point for social media service design.
However, beyond the above-mentioned Grön and Nelimarkka (2020), social media design research has not examined values through the lens of political ideologies. Instead, it seems that political ideologies form a design challenge in the context of social media service design: how to ensure that people across the ideological spectrum interact to create common ground (among many others Combs et al., 2023; Feltwell et al., 2020; Garimella et al., 2017). Alternatively, the broader social media research sees ideology as a background variable explaining human behaviour or content produced (for example, Morgan et al., 2013; Zhang and Counts, 2015).
Our brief review highlights the role of values in the debate on social media service design. It critiques the current situation and suggests alternative design approaches. At the same time, we highlight that, from this perspective and more widely in human–computer interaction research, these values are rarely understood in terms of political ideologies.
@@ 3.1. Research Setting
This study is situated in Finland, a Nordic country with 17 registered political parties, of which nine have representatives in the Finnish Parliament. Like in the rest of Europe, the media and political scientists use the two-dimensional ideological model (Figure 1) to categorise these parties (Grönlund and Söderlund, [n. d.]). In recent years, the ideological dimensions have become more prominent and affective polarisation has increased (Kekkonen et al., 2024).
In terms of political power, there are three equally large parties, each with about 20% of the popular vote in the most recent elections: National Coalition Party, the Finns Party, and Social Democratic Party. National Coalition Party and the Finns Party are political right, whereas the Social Democratic Party is political left. The Finns party is the only conservative party in this group. Following them, there are three somewhat smaller parties, with between 5% and 10% of the popular vote: the Center Party, the Greens Party, and the Left Alliance, The Greens party and the Left Alliance are left-leaning parties, while the Centre party is right-leaning and the only conservative party in this set of parties. All parties above have elected municipal, national and European representatives and have been part of the current or previous national government. The three remaining parties Swedish People’s Party of Finland Christian Democrats, and Movement Now, all had less than five per cent of the popular vote.
@@ 3.2. Participants
In total, we organised four workshops with 13 participants representing five111 We reached out to all parties, but even after several attempts, one party did not respond to our request. This also limited our ability to organise mixed and right-leaning workshops. of the six major parties. These parties included both right-leaning and left-leaning parties and included parties with more centralist stances and more extream stances. These parties represent the Finnish political spectrum broadly: together these five parties gained more than 60% of the popular vote, with the missing party adding an additional 20% of popular vote to the mix; thus our original target represented 86% of total popular vote. We contacted the party’s general secretary, executive director, or similar high-level officials. We asked them to suggest candidates for the workshop with expertise in social media and familiarity with the party’s political positions. Their suggestions included city council members, party office staff and youth organisation members, who we would consider to have a clear political ideology. Of the participants, six identified as women, six identified as men, and one identified in the other gender category. Their ages ranged from the early twenties to the seventies, with a median of 34 and a mean of 39. As Finland is a relatively small country, we do not provide more detailed demographic details to ensure the anonymity of participants.
The candidates suggested by party officials were suitable participants for this study. First, they have experience discussing politics through computer-mediated means and face-to-face settings. All participants discussed politics daily or often in face-to-face settings. Majority also discussed politics daily or often through instant messaging apps (5 daily, 6 often and 2 sometimes) and on social media (7 daily, 2 often and 4 sometimes). Second, they had an established ideological stance. All participants were able to indicate a closer political party and this party position was well established.
The groups had varied ideological compositions (see Table 1): one consisted of solely right-leaning participants (Group 1), two consisted of solely left-leaning participants (Groups 2 and 3), and one were mixed, including both right and left-leaning participants (Group 4). With the exception of Group 1222 Disappointingly, this group had a now-show from the other right-leaning party in our participant pool. This lead to the unfortunate situation, that this group was homogeneous in terms of political parties. As the logic of the workshops was complicated, we opted to run this group to ensure that we have at least one right-leaning group in our sample., all groups had participants from several parties. This group composition allowed us to examine the elicitation of political preferences in the left–right leaning axis, thus engaging how ideological thinking impacts research activities.
@@
Left (0)–Right (10) 8.0 1.33 3.50 5.50
@@
Conservative (0)–Liberal (10) 6.00 7.00 8.75 6.75

### v2

@@ Abstract.
Like any form of technology, social media services embed values. To examine how societal values may be present in these systems, we focus on exploring political ideology as a value system. We organised four co-design workshops with political representatives from five major parties in Finland to investigate what values they would incorporate into social media services. The participants were divided into one right-leaning group, two left-leaning groups, and one mixed group. This approach allows us to examine the differences in social media services designed by groups with different political ideologies i.e., value systems. We analysed produced artefacts (early-stage paper mockups) to identify different features and affordances for each group and then contrasted the ideological compositions. Our results revealed a clear distinction between groups: the right-leaning group favoured market-based visibility, while left-leaning groups rejected such design principles in favour of open profile work. Additionally, we found tentative differences in design outcomes along the liberal–conservative dimension. These findings underscore the importance of acknowledging existing political value systems in the design of social computing systems. They also highlight the need for further research to map out political ideologies in technology design.
@@
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW365publicationmonth: 11doi: 10.1145/3757546ccs: Human-centered computing Collaborative and social computingccs: Applied computing Law, social and behavioral sciencesccs: Social and professional topics Computing / technology policy
@@ 1. Introduction
Researchers have often focused on end-users and less on decision-makers, such as political elites, when considering different values and their influence on design (with the exception of Grön and Nelimarkka, 2020). At the same time, the government is a major provider of services, facilitated increasingly through digital means. Even more importantly, in democratic societies, the elected representatives have the legislative power – they establish laws increasingly related to technology, as shown in the European Union’s AI Act or the California Consumer Privacy Act. Both involve value-based decisions about the rights and responsibilities of citizens and their expected roles. In the political system, there are reasons to suggest that such decisions mirror a wider set of beliefs and assumptions (or political ideologies), as well as political realities based on different balances of power among representatives. Members of Congress, Parliament, or city councils, and other representatives rarely have been studied as stakeholders in the design of digital systems.
@@
We examined the outcomes of the workshops as the participants’ suggestions and analysed those towards the specific features and implicit affordances. Based on the literature on social media platforms and their affordances (Boyd, 2010; Bucher and Helmond, 2017; DeVito et al., 2017; Van Raemdonck and Pierson, 2021), we categorised affordances and connected them to underlying values to uncover ideological tensions between the different design solutions.
We next discuss relevant literature to present political ideologies as values and discuss social media design further. Following this, we describe the organisation of the co-design workshops. In this paper, we focus only on the final artefacts of the co-design activity, which exemplify the value-based decisions the groups made during the design workshop. We present a method to analyse these artefacts and summarise observations from the artefacts. We observe that in our sample, the right-leaning group supported market-based dynamics while left-leaning groups supported diverse self-presentation. The mixed group seemed to have both kinds of features. Before concluding our work, we discuss the importance of ideological diversity in design activities. Our work makes the following contributions focused on the design of information technology:
@@
  • We show how political ideologies (right–left) correspond with a particular type of social media design and identify design principles of flexible identity and competitive logic to be specific to the left- and right-leaning parties in our sample.
@@ 2. Background
Our work is based on the well-established observation that technology design and development are not value-neutral activities (among many others Nissenbaum, 2005; Winner, 1985). Winner (1985) infamously asked if artefacts have politics following a story about the bridges on Long Island, New York, USA. The bridges were constructed too low for public transportation buses to pass under them, thus removing access to the area for people using public transportation, primarily black people at that time. Winner argued that this is related to the designers’ social class bias and racial prejudice. Therefore, the bridges became manifestations of their designer’s political values. In the context of human–computer interaction, Keyes et al. (2019) has similarly called out the often apolitical nature of work, which they say to be implicitly neo-liberal in nature, calling out for more diverse ideological work.
@@ 2.1. Political Ideologies as a Set of Values
Values are conceptualised in different ways, and these conceptualisations often differ between academic disciplines.For example, political scientists often focus on political ideologies instead of values. Political ideologies are a set of values, beliefs, and attitudes people share that impact how they perceive and act in the world (Denzau and North, 1994). Instead of comprising single values, they formulate a value system. These political ideologies have practical consequences: within representative bodies, people with similar ideologies tend to form a political party (or a social movement in their early stages) to collaborate in advocating their values. The existence of cohesive parties allows one to map out ideologies’ dimensions; in the United States of America, this is often represented as a single dimension (republican–democrat) where various societal conflicts, such as economic stances, the role of the government, or racial assumptions, are positioned.
@@
Instead of a single ideological axis, a two-dimensional model is used across Europe (where our study is situated) (see Figure 1). The first axis corresponds to different values in terms of the degree of economic freedom (left–right), and the second axis corresponds to the level of desirable societal diversity (value conservative–liberal or green-alternative-liberal–traditional-authoritarian-nationalist) (Caughey et al., 2019; Kriesi et al., 2006). For example, a supporter or member of a left-liberal party, such as social democrats or the greens, might push for the right to self-define gender identity and extensive government oversight in the economic operation. In contrast, a supporter or member of a right-conservative party, such as Christian democrats or conservative parties, might push for traditional gender roles and limit the government’s role in the economic affairs. In general, while the model is simplistic, it has provided extensive value for empirical research on political decision-making, political opinions, and voting (e.g., Caughey et al., 2019; Huber and Inglehart, 1995; Hooghe et al., 2002; Knight, 2006). For example, political scientists have identified what kind of social and economic policies are commonly used across the two-dimensional model (Cochrane, 2015a; Jahn, 2011; Däubler and Benoit, 2022). In addition, some research has focused on the interconnection between technology and the party divisions: There is mixed evidence that left-leaning political parties are adapting information and communication technology more extensively to allow contact and participation in party functions (González-Cacheda and Cancela Outeda, 2024).
@@
As ideologies are systems based on a set of values, these values can manifest themselves in the design and use of digital services. Examining the design of social media services, Grön and Nelimarkka (2020) illustrate that the left–right axis relates to different degrees of paternalistic values: left-leaning participants supported a higher degree of paternalism while right-leaning participants highlighted the importance of individual responsibility. Their work highlights ideologies as an area of political competition, where individuals with different ideologies prefer design approaches aligned with their worldview. Relatedly, Détienne et al. (2019) built on the concept of ideologies to understand community memberships in conflict. For example, they identified an ideological conflict in the editing of Wikipedia articles. They observed strong disagreement between authors with a religious point of view and those with a scientific point of view, leading to an edit war and personal attacks on the platform. This instance underscores the potential for conflict and disagreement between political ideologies, making it clear that finding a compromise on an agreed solution can be difficult and sometimes impossible. This work highlights the importance of reflecting more expansive sets of values (or ideologies) and understanding them through wider groups, even if their ideologies do not fit the two-dimensional political ideology model presented above.
@@
The above examples illustrate the value of approaching not only individual values but sets of values, in other words, ideologies, in contrast to more prevalent studies focusing on specific values in context of information system design. The focus on political ideologies – pushed forward by political parties – provides the additional benefit of reflecting the legislative and executive processes at the societal level.
@@ 2.2. Values in the Context of Social Media Services
Like any sociotechnical arrangement, social media services incorporate and display values. The most prominent value tension in contemporary public discourse concerns social media companies’ for-profit operations and their contrast with other values, such as human rights. Thus, the discourse has a focus on capitalism. For example, Zuboff (2018) has argued that we are now in the era of surveillance capitalism, in which the collection and use of personal data has become a widespread business model. Social media companies use such personal data for targeted advertising, but while doing so, limit our right to privacy. An alternative line of criticism is the focus on user engagement and attention as a (capitalist) metric, leading to the promotion of affective content and thus having negative implications for political discourse on these platforms.
Given these concerns, there are proposals on alternative value bases for social media services. Jia et al. (2024) discuss how the news-feed algorithm could promote content more aligned with democratic ideals by sorting content so that it does not show support for undemocratic practices, partisan animosity or violence, or social distrust. Others have proposed different normative bases for rethinking social media services (for summary, see Goldberg, 2016). For example, the ideas that social media services could serve as a place for political deliberation have been repeated in the literature (e.g., Semaan et al., 2014, 2015). Similarly, researchers have shown that these values are culturally situated; what might be accepted in one cultural context may be rejected in another (e.g., Nelimarkka et al., 2019) and other researchers have called for more participatory and democratic approaches to governance (Fan and Zhang, 2020; Zhang et al., 2020). These examples again illustrate the critical role of values as a standing point for social media service design.
However, beyond the above-mentioned Grön and Nelimarkka (2020), social media design research has not examined values through the lens of political ideologies. Instead, it seems that political ideologies form a design challenge in the context of social media service design: how to ensure that people across the ideological spectrum interact to create common ground (e.g., Combs et al., 2023; Feltwell et al., 2020; Garimella et al., 2017; Nelimarkka et al., 2018). Alternatively, social media research sometimes positions ideology as a background variable that explains human behaviour or the content produced (for example, Morgan et al., 2013; Zhang and Counts, 2015a).
Our brief review above highlights the role of values in the debate on social media service design. It critiques the current situation and suggests alternative design approaches. At the same time, we highlight that, from this perspective and more widely in human–computer interaction research, these values are rarely understood in terms of political ideologies.
@@ 3.1. Research Setting
This study is situated in Finland, a Nordic country with a multiparty political system with 17 registered political parties, of which nine have representatives in the Finnish Parliament. Like in the rest of Europe, the media and political scientists use the two-dimensional ideological model (Figure 1) to categorise these parties (Grönlund and Söderlund, 2024). In recent years, the ideological dimensions have become more prominent and affective polarisation has increased (Kekkonen et al., 2024).
In terms of political power, there are three equally large parties, each with about 20% of the popular vote in the most recent elections: National Coalition Party, the Finns Party, and Social Democratic Party. Following Grönlund and Söderlund (2024), National Coalition Party and the Finns Party are political right, whereas the Social Democratic Party is economic left. The Finns party is the only value-conservative party in this group, while both National Coalition Party and the Social Democratic party belong to the value-liberal side of parties. Following them, there are three somewhat smaller parties, with between 5% and 10% of the popular vote: the Center Party, the Greens Party, and the Left Alliance, The Greens party and the Left Alliance are left-leaning and value-liberal parties, while the Centre party is right-leaning and the only value-conservative party in this set of parties. All parties above have elected municipal, national and European parliament representatives and have been part of the current or previous national government. The three remaining parties Swedish People’s Party of Finland, Christian Democrats, and Movement Now, all had less than five per cent of the popular vote. Ideologically, the first one belongs to the economic right and value-liberal quadrant, while the second belongs to economic right and value-conservative quadrant; the sample is too small to make definitive position for the Movement now, but within popular discussion, it is a value-liberal economic right party.
When positioned on the two-dimensional model of political ideologies (Figure 1), the parties are situated in the value-liberal economic left, value-liberal economic right, and value-conservative economic right quadrants. Beyond these, there are various specific issues for each party; for example, the Swedish People’s Party of Finland strongly supports Finland as bilingual country (with both Finnish and Swedish serving as official languages). While these specific issues are not captured by the two-dimensional model, the two-dimensional model is often used to describe the ideological landscape in Finland (e.g., Grönlund and Söderlund, 2024; Kekkonen et al., 2024; Ford and Jennings, 2020b; Huber and Inglehart, 1995).
@@ 3.2. Participants
In total, we organised four workshops with 13 participants representing five111 We reached out to the six largest parties, but even after several attempts, one party did not respond to our request. This also limited our ability to organise mixed and right-leaning workshops. of the six major parties. These parties included both right-leaning and left-leaning parties and included parties with more centralist stances and more extreme stances. These parties represent the Finnish political spectrum broadly: together these five parties gained more than 60% of the popular vote, with the missing party adding an additional 20% of popular vote to the mix; thus our original target represented 86% of total popular vote. We contacted the party’s general secretary, executive director, or similar high-level officials. We asked them to suggest candidates for the workshop with expertise in social media and familiarity with the party’s political positions. Their suggestions included city council members, party office staff and youth organisation members, who we would consider to have a clear political ideology (we further confirmed this with a survey, as indicated later). Of the participants, six identified as women, six identified as men, and one identified in the other gender category. Their ages ranged from the early twenties to the seventies, with a median of 34 and a mean of 39. As Finland is a relatively small country, we do not provide more detailed demographic details to ensure the anonymity of participants.
@@
The candidates suggested by party officials were suitable participants for this study. First, they have experience discussing politics through computer-mediated means and face-to-face settings. All participants discussed politics daily or often in face-to-face settings. The majority also discussed politics daily or often through instant messaging apps (5 daily, 6 often and 2 sometimes) and on social media (7 daily, 2 often and 4 sometimes). Second, they had an established ideological stance. All participants had a well-established political party identity, i.e. there was a party they felt closest to. In addition, we conducted a survey of political ideologies on the economic left—right and value conservative–liberal axis222 We used the survey items used in the most recent Finnish National Election Survey to evaluate ideological positions and discussed with the project lead of the National Election Survey on the suitability of these indicators. to confirm that these individuals self-described their political ideology similar to the ideology position of their party (see Figure 2).
@@
Economic left (0)– Economic right (10) 8.0 1.33 3.50 5.50
@@
Value conservative (0)– Value liberal (10) 6.00 7.00 8.75 6.75
@@
The groups had varied self-reported ideological compositions (see Table 1): one consisted of participants solely right-leaning (Group 1), two consisted of participants solely left-leaning (Groups 2 and 3), and one was mixed, including participants both right- and left-leaning (Group 4). With the exception of Group 1333Unfortunately, this group had a no-show from the other right-leaning party in our participant pool. This led to the unfortunate situation that this group was homogeneous in terms of political parties. As the logistics of the workshops were complicated, we nevertheless opted to run this group to ensure that we have at least one right-leaning group in our sample., all groups had participants from several parties. This group composition allowed us to examine the elicitation of political preferences in the left–right leaning axis, thus engaging how ideological thinking impacts research activities.